OpenStudy (anonymous):

solve : ^3squareroot x-7 =2

OpenStudy (anonymous):

so its \[\sqrt[3]{x-7} = 2\]

OpenStudy (anonymous):

to get rid of the cube root sign cube both sides of the equation than solve for x \[x -7 = 2^{3}\] , solve for x?

OpenStudy (anonymous):

x=15 ?

OpenStudy (anonymous):

yes good.